PLEASE HELP!!!!!!!!!!!
In a recent poll of 1,600 homeowners in the United States, one in eight homeowners reports having a home equity loan that he or she is currently paying off. Using a confidence coefficient of 0.99, derive the interval estimate for the proportion of all homeowners in the United States that hold a home equity loan. (You may find it useful to reference the z table. Round intermediate calculations to at least 4 decimal places. Round "z" value and final answers to 3 decimal places.)
Thank you very much for your help!!

Answer by Boreal(15235)

half-interval for a 99% CI is z(0.995)*sqrt(p*(1-p)/n. p=1/8 1-p=7/8 and their product is 1/64
sqrt term is sqrt(7/64)/sqrt(1600)=sqrt(7)/8*40

=2.576*(sqrt(7/64))/40
=2.576*sqrt(7)/320
=0.0213
interval is 1/8+/-0.0213
(0.1037, 0.1463) or (0.104, 0.146)